OpenStudy (anonymous):

A highway makes an angle of 6 degrees w/ the horizontal. This angle is maintained for a horizontal distance of 5 miles. To the nearest hundredth of a mile, how high does the highway rise in this 5 mile section? Show steps

OpenStudy (anonymous):

I was thinking 

OpenStudy (anonymous):

tan 6 degrees = x/5 Am I on the right track ?

OpenStudy (paki):

you know here "theta" is 6 degrees... so using the relation... tan(6) = rise/5 rise = 5*tan(6) ... solve now....

OpenStudy (paki):

yup your procedure is Also rite...

OpenStudy (paki):

what you got the value of rise...

OpenStudy (anonymous):

tan 6 = .105

OpenStudy (anonymous):

5*tan 6 = .526

OpenStudy (paki):

how you got .105 and .526...?

OpenStudy (paki):

just do like this.... tan (6) = rise/5... rise = 5*tan(6)... so rise = 0.525 miles

OpenStudy (paki):

@Shay17 now got it...

OpenStudy (anonymous):

Ok .. i'm still confused

OpenStudy (paki):

where....? i hope the calculations are all simple and easy...

OpenStudy (anonymous):

They seem to be but I'm still confused on how I messed up

OpenStudy (paki):

where you are confused... tell me...?

OpenStudy (anonymous):

Idk how did I mess up I just wanna make sure I fully understand

OpenStudy (paki):

ahhh. so then have a look on my calculation and hope you will be satisfied... :)

OpenStudy (anonymous):

Yes so what would a good explanation be ?

OpenStudy (anonymous):

So wait the highway rises .53 miles in the 5 mile section ?

OpenStudy (paki):

yes... in 5 miles it has a slope of 0.53 miles...

OpenStudy (anonymous):

I'm good thank you I just had to re read it again ! :)

OpenStudy (paki):

ahhh pleasure :)

OpenStudy (anonymous):

@iambatman